OpenStudy (anonymous):

There is an inverse relationship between x and y. If y is twenty when x is 2, what is x when y is 8?

OpenStudy (anonymous):

Let the relationship is y = k/x. Put y = 20 and x= 2 and find k

OpenStudy (anonymous):

i dont even know tbh

OpenStudy (neer2890):

ok. so we have inverse relationship between x and y means when x increases , y decreases and when y increases ,x decreases. now 

OpenStudy (neer2890):

we use chain rule to solve this. if we found inverse relationship between two quantities then we reciprocal them so, \[\frac{ 2 }{ ? }=\frac{ 8 }{ 20 }\] now solve for ? which is the required value.

OpenStudy (neer2890):

\[?=\frac{ 20*2 }{ 8 }=5\]

OpenStudy (anonymous):

thank you!!!!! :') @neer2890